OpenStudy (anonymous):

please help me in algebra 1!! Task 2 Part 1. Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model. a√x+b+c=d Use a constant in place of each variable a, b, c, and d. You can use positive and negative constants in your equation. Part 2. Show your work in solving the equation. Include the work to check your solution and show that your solution is extraneous. Part 3. Explain why the first equation has an extraneous solution and the second does not.

OpenStudy (anonymous):

any particular question ?

OpenStudy (anonymous):

i dont even know what it wants me to do

OpenStudy (anonymous):

please i at least need solution to part one!

OpenStudy (anonymous):

\[a\sqrt{x+b}+c=d\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and you want it to have a solution? lets work backwards

OpenStudy (anonymous):

well i need to make an equation that is extraneous and not extraneous

OpenStudy (anonymous):

lets put \[2\sqrt{x+1}+2=6\] and solve \[2\sqrt{x+1}+2=6\\ 2\sqrt{x+1}=4\\ \sqrt{x+1}=2\\x+1=4\\ x=3\] in this case \(3\) is not an extraneous solution because it works

OpenStudy (anonymous):

ok so how do you make it extraneous?

OpenStudy (anonymous):

you could make the number on the right negative that would work try \[2\sqrt{x+1}+2=-1\]

OpenStudy (anonymous):

or to make it easier try \[2\sqrt{x+1}+2=-4\]

OpenStudy (anonymous):

whatis extraneous though?

OpenStudy (anonymous):

solve and see that it is "extraneous" in that the solution does not work

OpenStudy (anonymous):

\[2\sqrt{x+1}+2=-4\\ 2\sqrt{x+1}=-6\\ \sqrt{x+1}=-3\\ x+1=9\\ x=8\] but \(8\) is not a solution to the original equation

OpenStudy (anonymous):

so its not extraneous?